OpenStudy (anonymous):

What is the vector projection of 's' onto 'r'? r=3i+4j-8k & s=2i-3j+7k

OpenStudy (zarkon):

use \[\frac{r\cdot s}{r\cdot r}r\]

OpenStudy (anonymous):

Find r-unit vector and the dot product of s and r-unit vector will result in the required projection